OpenStudy (anonymous):

A.Write the expression (-64)^2/3 in radical notation. B. Evaluate the radical expression. A. Write the expression in radical notation. Do not evaluate. (-64)^2/3=

OpenStudy (anonymous):

n-e body? lol.

OpenStudy (anonymous):

\[A.\text{ }(-64)^{2/3}=\sqrt[3]{-64^2} \] \[B.\text{ }(-64)^{2/3}==16 \]

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

Hmmmm.... it wasn't right. :(

OpenStudy (anonymous):

Am I missing something?

OpenStudy (anonymous):

It does say to type an exact answer, using radicals as needed

OpenStudy (anonymous):

theycallmekelly, Do you want a comment from moi?

OpenStudy (anonymous):

please.......

OpenStudy (anonymous):

DO you know the correct answer?

OpenStudy (anonymous):

...... I'm so sleepy.....

OpenStudy (anonymous):

You quit writing robtobey!

OpenStudy (anonymous):

With regard to problem A. I took it to mean that the fractional exponent should be converted to radical form, ie: radical sign or square root symbol. Problem B seems to request the value of Problem A. \[(-64)^2 = 4096\] The cube root of 4096 is 16. \[4096^{1/3} = \sqrt[3]{4096} = 16\] By convention: \[\sqrt[3]{x^2} = x^{\frac{2}{3}} \]

OpenStudy (anonymous):

So A would be 4096?

OpenStudy (anonymous):

...........

OpenStudy (anonymous):

No. Refer to the second version of A. It says not to evaluate A. Only show the symbolic form of the expression, not the number value of the expression. The whole thing seems to be a test of whether or not you under stand how to handle fractional exponents. Google "fractional exponents math' and I'm sure you will find some excellent presentations regarding fractional exponents.

OpenStudy (anonymous):

Lol, so what is A then?

OpenStudy (anonymous):

The answer to A is: \[\sqrt[3]{(-64)^2}\] That's about all I can do with this problem.

OpenStudy (anonymous):

Yes! Lol, that part was right!

OpenStudy (anonymous):

Any idea on B?

OpenStudy (anonymous):

The answer is 16 as I tried to explain in my second response to you tonight.